### v2

@@ 6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.32 admits a unique adapted solution. Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.30 guarantees the existence and uniqueness of solution. Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.32) with the same formulation and the same $(\rm H_{0})$} condition in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
@@
  • $(\rm H_{q}^{\prime})$}
    There exist constants $L_{z}>0$, and $K_{z}>0$ such that:
@@
This is the fully coupled quadratic case, in contrast with the separately quadratic $(\rm H_{q})$} of the previous sections. We provide an explicit example here.
Example 6.5.
For simplicity, consider the one-dimensional case $d=1$ and define
This $g$ satisfies the fully coupled quadratic condition $(\rm H_{q}^{\prime})$}, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
First, we note that the terminal condition in BSDE 6.30 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.32 in terms of the centered random variable $\eta$:
@@
$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi]$
are both well-defined because of $(\rm H_{0})$}.
@@
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
@@
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}}.$
Then the BSDE 6.33 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
@@
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T.$
@@
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s}.$
@@
$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right]^{2}.$ (6.34)
Together with the $(\rm H_{q}^{\prime})$} condition and Jensen’s inequality, we have
$\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{z}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
@@
$\mathbb{E}\left[\int_{t}^{T}\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|ds\mid\mathcal{F}_{t}\right]$ (6.35)
$\leq\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right].$
@@
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]\right]^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]^{2}\right)$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{z}^{2}\left\|z\right\|_{BMO}^{4}\right).$
The last inequality uses Jensen’s inequality:
$\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right).$
Now, using $a^{2}+b^{2}\leq(|a|+|b|)^{2}$, it follows that
@@
where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{z}$. To obtain a uniform bound, set $R>0$ such that
@@
is then such that $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
@@
$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\mid\mathcal{F}_{t}\right)\right]^{2}$
$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$
$\quad+16K_{z}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\mid\mathcal{F}_{t}\right]\mathbb{E}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\mid\mathcal{F}_{t}\right]$
$\leq\left(8L_{z}^{2}T+32K_{z}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
@@
$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{z}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
@@
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
@@
Let $\delta:=(65L_{z}^{2})^{-1}<(64L_{z}^{2})^{-1}$, and divide $[0,T]$ into $N$ subintervals of length at most $\delta$: $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{k+1}-t_{k}\leq\delta$ for all $k$. By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{z}}=\frac{1}{65^{2}K_{z}}$. We construct the solution recursively, starting from the last interval $t_{N-1},t_{N}]$. On this interval, the terminal value is $\eta$, and the smallness condition is satisfied by assumption. Thus, by applying the local existence and uniqueness lemma, we obtain a unique solution $(U,Z)$ on $t_{N-1},t_{N}]$ with a priori bound in $\mathcal{B}_{\rho}$. Then we proceed recursively: at each interval $t_{k-1},t_{k}]$, we solve the BSDE using $U_{t_{k}}$ as the terminal value. The a priori bound $\|U\|_{\mathbb{S}^{\infty}}\leq\rho$ ensures that the smallness condition $\|U_{t_{k}}\|_{L^{\infty}}\leq\rho$ continues to hold at each step. Finally, By concatenating these local solutions, we obtain a unique adapted solution $(U,Z)$ on $[0,T]$. ∎
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.33, the solution to the original BSDE 6.32 is obtained by the explicit transformation
$Y_{t}=U_{t}+g(t,0,0)+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
@@
First we will show the driver $f$ (6.31) satisfies the condition $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. The verification of the first condition is straightforward due to the boundedness of $\theta$. To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
@@
Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\rm H_{q}^{\prime})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.30) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
@@
is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above. This verification follows directly from arguments in Theorem 7 of hu2005utility, and we omit the details here. ∎

### v3

@@ 6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first adapt ideas from Tevzadze (2008): under a smallness condition on the centred terminal value and a (local) Lipschitz component, the fully coupled mean-field BSDE (6.33) admits a unique solution. We then globalize this local result by a localization/concatenation procedure in Lemma 6.6 and Proposition 6.7. Finally, we return to the utility maximization problem and identify both the value function and an optimal strategy in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.33) with the same formulation and the same $(\mathrm{H}_{0})$ condition in the main result but with a different representation and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
@@
  • $(\mathrm{H}_{q}^{\mathrm{LK}})$}
    There exist constants $L_{z}>0$, and $K_{q}>0$ such that:
@@
This is again the fully coupled quadratic case, that is, $(\mathrm{H}_{q}^{\mathrm{unif}})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$} are equivalent up to a change of constants. We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
First, we note that the terminal condition in BSDE 6.31 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.33 in terms of the centered random variable $\eta$:
@@
$U_{t}:=Y_{t}-\int_{0}^{t}g(s,0,0)ds-\mathbb{E}[\xi]\text{ and }\eta:=\xi-\mathbb{E}[\xi]$
are both well-defined because of $(\mathrm{H}_{0})$.
Remark 6.5.
When $b$ and $\sigma$ are deterministic, (6.34) has a unique solution if we assume the terminal condition $\xi$ satisfies the condition described in Theorem 2.6.
In the following, we give the results when $b$ and $\sigma$ are not deterministic. In this case, we need smallness in the spread of the terminal condition from its mean.
@@
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$ and terminal condition $\eta$ be a bounded terminal condition such that
@@
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{q}}.$
Then the BSDE 6.34 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
@@
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T.$
@@
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s}.$
@@
$|U_{t}|^{2}+\mathbb{E}_{t}\left[\int_{t}^{T}|Z_{s}|^{2}ds\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}_{t}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\right]^{2}.$ (6.35)
Together with the $(\mathrm{H}_{q}^{\mathrm{LK}})$} condition and Jensen’s inequality, we have
$\left|\int_{\tilde{\Omega}}\big[g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{q}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
@@
$\mathbb{E}_{t}\left[\int_{t}^{T}\big|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\big|ds\right]$ (6.36)
$\leq\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right].$
@@
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]\right]^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]^{2}\right)$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{q}^{2}\left\|z\right\|_{BMO}^{4}\right).$
The last inequality uses Jensen’s inequality: $\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right)$. Now, using $a^{2}+b^{2}\leq(|a|+|b|)^{2}$, it follows that
@@
where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{q}$. To obtain a uniform bound, set $R>0$ such that
@@
satisfies $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
@@
$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}_{t}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\right)\right]^{2}$
$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}_{t}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\right]$
$\quad+16K_{q}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}_{t}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\right]\mathbb{E}_{t}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\right]$
$\leq\left(8L_{z}^{2}T+32K_{q}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
@@
$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{q}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
@@
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$. Suppose there exists $\rho>0$ such that
@@
Let $\delta:=(65L_{z}^{2})^{-1}<(64L_{z}^{2})^{-1}$, and divide $[0,T]$ into $N$ subintervals of length at most $\delta$: $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{k+1}-t_{k}\leq\delta$ for all $k$. By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{q}}=\frac{1}{65^{2}K_{q}}$. We construct the solution recursively, starting from the last interval $t_{N-1},t_{N}]$. On this interval, the terminal value is $\eta$, and the smallness condition is satisfied by assumption. Thus, by applying the local existence and uniqueness lemma, we obtain a unique solution $(U,Z)$ on $t_{N-1},t_{N}]$ with a priori bound in $\mathcal{B}_{\rho}$. Then we proceed recursively: at each interval $t_{k-1},t_{k}]$, we solve the BSDE using $U_{t_{k}}$ as the terminal value. The a priori bound $\|U\|_{\mathbb{S}^{\infty}}\leq\rho$ ensures that the smallness condition $\|U_{t_{k}}\|_{L^{\infty}}\leq\rho$ continues to hold at each step. Finally, By concatenating these local solutions, we obtain a unique adapted solution $(U,Z)$ on $[0,T]$. ∎
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.34, the solution to the original BSDE 6.33 is obtained by the explicit transformation
$Y_{t}=U_{t}+\int_{0}^{t}g(s,0,0)ds+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
@@
First we will show the driver $f$ (6.32) satisfies the condition $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$}. The verification of the first condition is straightforward due to the boundedness of $\theta$. To verify $(\mathrm{H}_{q}^{\mathrm{sep}})$}, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
@@
Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\mathrm{H}_{q}^{\mathrm{LK}})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.31) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
@@
is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above. This verification follows directly from arguments in Theorem 7 of Hu et al. (2005), and we omit the details here. ∎